### v1

Abstract
Research in various fields is currently experiencing challenges regarding the reproducibility of results. This problem is also prevalent in machine learning (ML) research. The issue arises primarily due to unpublished data and/or source code and the sensitivity of ML training conditions. Although different solutions have been proposed to address this issue, such as using ML platforms, the level of reproducibility in ML-driven research remains unsatisfactory. Therefore, in this article, we discuss the reproducibility of ML-driven research with three main aims: (i) identify the barriers to reproducibility when applying ML in research as well as categorize the barriers to different types of reproducibility (description, code, data, and experiment reproducibility), (ii) identify potential drivers such as tools, practices, and interventions that support ML reproducibility as well as distinguish between technology-driven drivers, procedural drivers, and drivers related to awareness and education, and (iii) map the drivers to the barriers. With this work, we hope to provide insights and contribute to the decision-making process regarding the adoption of different solutions to support ML reproducibility.
Keywords:
Machine Learning Artificial Intelligence Reproducibility Irreproducibility
1 Introduction
As in many scientific fields (e.g., [8, 66]), research in artificial intelligence (AI) in general, and machine learning (ML) in particular, faces crucial doubts over the reproducibility of research [35]. This has raised concerns about the reliability and validity of many scientific findings, risking diminishing confidence in the overall body of scientific knowledge. Misleading or irreproducible results can lead to wasted resources, hinder scientific progress, reduce trust in science, and impact decision-making in various fields [49, 28]. The “reproducibility crisis” has been a topic of discussion since the mid-2010s, particularly in the medical and behavioral science fields [8, 53]. Recently, this issue has also been associated with research involving ML methods [35]. Apart from the common challenges faced by many disciplines, such as limited sharing of data and source code, the use of ML introduces unique challenges for reproducibility, including sensitivity to ML training conditions, sources of randomness [63], and the increasing use of AutoML tools [41].
While the literature on these issues is expanding [45, 29, 32, 39, 25, 4, 68, 26], there is still no comprehensive overview of the associated barriers and drivers. For example, in [26], the authors identify and categorize sources of irreproducibility in ML and how these sources affect conclusions drawn from ML experiments. However, this study does not investigate drivers to address these sources of irreproducibility. Thus, our paper provides a contextual categorization of the barriers and drivers to the four types of ML reproducibility proposed by [25], with specific reference to research in both Computer Science and Biomedical fields. We also propose a Drivers-Barriers-Matrix to summarize and visualize the results of the discussion. Such an analysis stands to clarify the current state regarding ML reproducibility, provide guidance for those new to the subject, and highlight potential gaps, solutions, and avenues for future exploration.
The paper is structured as follows: in Section 2, we clarify terms and working definitions. We then analyze the barriers to increased reproducibility of ML-driven research (Section 3), and next the drivers that support ML reproducibility, including different tools, practices, and interventions (Section 4). Finally, we map the barriers to the drivers to help determine the feasibility of various options for enhancing ML reproducibility (Section 5). We close the paper with a conclusion and an outlook into our future research in Section 6.
2 Defining Reproducibility
The concept of reproducibility can have different interpretations across various research fields and even within the same field [22]. To address this confusion, reproducibility is broadly defined and then further categorized into various types or degrees. A key distinction is made by Goodman et al. [24] who specify a fundamental division between whether we (i) mean reproducible in principle (termed “methods” reproducibility) due to sufficient description/sharing of methodologies, materials, etc., or (ii) whether results/conclusions actually prove to be reproducible when experiments or analyses are re-done. In the second category, they distinguish “results” and “inferential” reproducibility, depending on whether it is the analyses or inferences to broader conclusions that are reproduced.
Within AI research, widely accepted definitions that build further on these key distinctions have been proposed by  [29, 25]. We follow and build upon these latter definitions, and hence here outline them in some length.
Gundersen et al. [27] define reproducibility in general as “the ability of independent investigators to draw the same conclusions from an experiment by following the documentation shared by the original investigators”.
Relating to point (ii) of Goodman et al.’s schema, Gundersen and colleagues  [29, 25] further distinguish the targets of reproducibility, i.e., how closely an experiment can be reproduced:
  • Outcome reproducibility requires the reproduced experiment to have the same or adequately similar outcome as the original experiment. Due to this, the same analysis and interpretation follow, and the hypothesis is either supported or rejected by both experiments.
  • Analysis reproducibility does not require the reproduced experiment to have the same/similar outcome; however, if the same/similar analysis and, therefore also, interpretation can be made, an experiment is analysis reproducible.
  • Interpretation reproducibility does not require the reproduced experiment to have the same/similar outcome nor analysis but requires the interpretation to be the same as the original one. In such a case, an experiment is interpretation reproducible.
Figure 1: Types of reproducibility. Adapted from Gundersen [25].
This categorization aims to overcome the problem of ambiguity when making specific claims about the reproducibility of an experiment. Often in literature, authors write about reproducing the same “results” of an experiment. It is not apparent, however, in which cases they mean to achieve the same computational outcome, i.e., outputs of the algorithms, or whether they mean to reach the same analysis or interpretation. Therefore, achieving interpretation reproducibility is a more general and often less stringent requirement than achieving outcome reproducibility. This categorization is not specific to ML but generally applicable to any research field that conducts computational experiments.
In addition, relating to point (i) of the Goodman et al. schema (“methods” reproducibility), [25] specifies degrees to which methods can be made transparent through description or sharing. These four types are defined as R1 Description, R2 Code, R3 Data, and R4 Experiment. The lower the level of reproducibility, the less shared information there is, making the experiment more difficult to reproduce. Giving an example: In general, all published research experiments are accompanied by a textual description of the experiment. If this textual description is the only information shared by the authors, the research is categorized as R1 Description, which, according to this scheme, is the lowest type of reproducibility. In contrast, if all three building blocks, i.e., text, code, and data, are shared, the experiment can be categorized as R4 Experiment, which is the highest type of reproducibility. Furthermore, the distinction between R2 Code and R3 Data is defined by the textual description of the original experiment that is accompanied by either code or data, respectively.
Also, the different types of ML reproducibility exhibit an interplay between generalizability and reproducibility. R1 Description leads to the highest generalizability but to the weakest reproducibility, while R4 Experiment leads to the highest reproducibility and lowest generalizability.
This relationship is illustrated in Figure 1 [25]. Furthermore, Figure 1 visualizes the building blocks associated with the different types of reproducibility. In what follows, unless stated otherwise, we use the definitions proposed by Gundersen and colleagues.
3 Barriers in ML Reproducibility
In the following, we discuss nine barriers to ML reproducibility, categorized into the four types of reproducibility mentioned in the previous section. Where applicable, we give examples within the research fields of Biomedical Science, and Computer Science.
3.1 R1 Description
3.1.1 Completeness and quality of reporting.
Research often lacks reproducibility due to missing or vague methodological details. Mainly, there are three issues in this regard, which often hinder the reproduction of study results [57]:
  1. 1.
    The ML model or training procedure is either incorrectly specified or under-specified. Reports should give clear details on all steps of the procedure, even if data and code are not shared. This includes details about which ML models are used, as well as details on the training data and data preprocessing.
  2. 2.
    The evaluation metrics used to report results are not properly specified. There are many metrics, which can be used to evaluate ML models, e.g., accuracy, receiver-operator-curve (ROC), or mean-squared-error (MSE). It is important to define these metrics and also explain why they were used.
  3. 3.
    Often results are selectively reported, e.g., researchers may only provide results for the best performance out of many test runs, instead of properly assessing and reporting the performance over multiple runs. For example, in research on Natural Language Processing (NLP), often, reproductions turn out to perform worse than stated in publications [10].
Generally, it is important for studies to use a robust methodology and provide detailed reports, for other researchers to verify results and understand how analyses were conducted. While ML models have proven highly effective in Biomedical fields, studies often fall short in providing comprehensive and high-quality reporting. For example, in studies of predicting cardiometabolic risk from dietary patterns [54] or supporting the clinical management of diabetes [38], ML models were observed to be very promising. While this further enhances the promise of applying ML models for various clinical prediction tasks, there is a clear need for thorough reporting and validation of these models to allow for their integration into routine clinical care [38]. This is also true for the application of ML models for cancer imaging, where ML models often surpass radiologists in performance, but publications on these models lack the details needed to reproduce the results [59].
3.1.2 Spin practices and publication bias.
Another flaw, which has been observed in ML-based research and negatively affects reproducibility, is “spin”. It refers to the misuse of language to “intentionally or unintentionally affect the interpretation of study findings”. Accordingly, it is also understood as an inconsistency between the study results and the conclusions, in the sense that results are over-generalized or the claimed conclusions are not supported by the scientific method. This has been shown to impact both the interpretations and decision-making by readers [5]. In ML-based Biomedical research, the most common practice of spin includes recommending models for various applications without providing external validation in the same study. More concretely, the recommendation to use a model either in a clinical setting or for a different population is only validated in approximately 15% of the cases. Other observed instances of spin are invalid comparisons of results to previous developments and the use of leading words and strong statements to make the results sound more prolific [5]. The use of spin in numerous publications might be partly attributed to the reward system in academia, which often values spin practices in research, as more prolific results are more likely to lead to accepted publications. Publication bias in general is a big problem in a variety of different regards. It seems that research from high-profile institutions is published more easily and subject to less scrutiny in review and there also seems to be a tendency to prefer more complex solutions over simpler ones, which solve the same problem [72].
3.2 R2 Code
3.2.1 Limited access to code.
Published ML research is often not accompanied by available data and code. Only one-third of researchers share data, and even fewer share their source code [36]. This can be for many reasons, such as the increasing pressure on researchers to publish quickly, which in turn does not allow them to polish the code and decreases the willingness to release the code, or simply due to private source code that builds on unpublished code. Sharing source code correctly requires publishing seven pieces of information: hypothesis, prediction method, source code, hardware specifications, software dependencies, experiment setup, and experiment source code. Only if the source code is accompanied by all these artifacts, it is specified sufficiently to allow reproducing an experiment [28]. Unfortunately, current research often does not meet this requirement, leading to reproducibility issues due to different software versions, hyperparameter settings, or hardware differences  [33, 10]. For example, in Computer Science, the lack of shared code and data significantly contributes to a lack of reproducibility [18]. Even when code and data are shared, they are often poorly documented, or the code is only provided as a skeleton instead of a fully executable code, which does not ensure reproducibility. Thus, shared code should also include the code for data preprocessing, hyperparameter tuning, and managing random seeds, as well as code for comparisons against baseline models.
3.3 R3 Data
3.3.1 Limited access to data.
The main barrier concerning R3 Data is that data is often not shared. This could be due to privacy concerns or a lack of incentives and motivation. Moreover, many benchmark and training datasets encounter challenges related to copyright, licensing, and longevity. These datasets may also raise ethical concerns, such as the unintentional inclusion of privacy-sensitive or harmful content, making it difficult to share the data for ML model training [55]. Similar to the sharing of source code, sharing only the datasets is insufficient. For proper use, it is also important to share specific splits, i.e., the training dataset, validation dataset, and test dataset [28]. Furthermore, data sharing needs to be accompanied by details about the provenance and preprocessing of data. This is crucial to investigate methodological errors such as data leakage and bias, which are described in the following.
3.3.2 Data leakage.
In practice, methodological issues such as data leakage often hinder the reproducibility of ML-based research [39]. This is due to the growing number of non-experts employing machine learning across different research fields [23], which is fueled by the ease of application of ML libraries and no-code off-the-shelf AI tools. In essence, data leakage happens when data on which the ML model should not be trained leaks into the training process. Data leakage can be categorized into 3 subcategories [39]:
  1. 1.
    No clean train/test split. Here, four variants are possible: (1) training data and test data are not split at all, (2) test data is also used to select the best features from the training data (feature selection), (3) test data is also used for imputation of missing data during preprocessing, and (4) duplicates occur in the training and test data.
  2. 2.
    Use of non-legitimate data. For example, when the use of antihypertensive drugs is used to predict hypertension. This data is non-legitimate since it would not be available in a real-world scenario and cannot realistically be used to predict hypertension for a new patient.
  3. 3.
    Test set is not drawn from the distribution of scientific interest. There are three possible variants: (1) temporal leakage, which is problematic for ML models that attempt to predict future outcomes, i.e., when some training samples have a later timestamp than samples available in the test set. (2) the training and test data are not independent of each other, e.g., there should not be samples in the training and test data which are drawn from the same person. (3) the test set is not chosen selectively; for instance, if the model is solely evaluated on data where it performs well.
3.3.3 Bias.
Bias in ML refers to the error introduced by approximating a real-world problem, which may be complex, by a simplified model, often leading to systematic deviations from the ground truth. Furthermore, this bias can arise when the model contains imbalances or reflects existing societal biases. ML models, which are subject to bias, do not generalize well and are therefore potentially problematic for ML reproducibility. There are eight bias variants that can arise during the data handling phase of ML development [65]: (1) selection bias - using data not being representative of the target group, (2) exclusion bias - excluding particular data samples based on the belief that they are unimportant, (3) measurement bias - favoring certain measurement results, (4) recall bias - labeling similar data samples differently, (5) survey bias - introducing data issues stemming from data collection surveys, (6) confirmation bias - favoring information, which confirms previous beliefs, (7) prejudice bias - including human-related prejudices in training data, and (8) algorithmic bias - replicating or amplifying biases by algorithms themselves. Biases, such as selection bias, often lead to the issue of validity shrinkage in Biomedical science research [37]. For example, in obesity and nutritional research, ML is used to predict obesity, heart rate, or the risk of heart attack, based on data of an individual. Here, validity shrinkage refers to the issue that a predictive model trained on a subset of data will most probably not perform well on new samples. The difference between predictive performance on known data and new data, however, is most often not accounted for in nutritional science and therefore also leads to performance claims that cannot be reproduced [37].
3.4 R4 Experiment
3.4.1 Inherent nondeterminism.
Inherent nondeterminism in ML models means that results can vary between test runs, even with identical code, data, and hyperparameters. This variation occurs from unanticipated influences during training, such as random parameter initialization, stochastic optimization, and random data subsampling (e.g., in k-fold cross-validation) [63, 43]. Neural networks are especially known for their inherent nondeterminism, leading to varied computational outcomes in multiple reruns due to numerous sources of randomness during training [1]. In some cases, inherent non-determinism can cause such large variations that reruns not only yield slightly different outcomes but also lead to significant fluctuations in the performance of an ML model [1]. This issue is exacerbated when other sources of variation, such as different hyperparameters, are introduced to the ML model. In such cases, the impact can be magnified, and it is often observed that minor changes in hyperparameters can result in significant performance loss [10]. For example, a review of the status quo of reproducibility in NLP research highlights some core issues with non-determinism that are exemplary for ML research [10]. Simply rerunning the original code during a reproduction can lead to different computational outcomes as the original reported ones. Reinforcement learning (RL), a subfield of ML, is particularly susceptible to these reproducibility issues, partially because of additional sources of non-determinism in the learning process, such as the RL environment or policy [50].
3.4.2 Environmental differences.
Various studies have demonstrated that hardware differences, such as different GPUs or CPUs, and compiler settings can lead to different computational outcomes [33]. Additionally, a comparison between the same ML algorithm with fixed random seeds executed using PyTorch [^1] https://pytorch.org/ and TensorFlow [^2] https://www.tensorflow.org/ resulted in different performances [58]. A comparison of the results of an ML experiment between different ML platforms also shows that out-of-the-box reproducibility is not provided. In general, it is questionable to what extent reproducibility can be ensured out-of-the-box [29]. Another important factor is the use of GPUs, which can increase randomness compared to the use of CPUs. This is due to parallel optimization and the use of optimizers in deep learning frameworks, such as PyTorch and TensorFlow. As a result, some researchers have resorted to solely using CPUs for computation. However, this comes at the expense of time-efficiency [3].
3.4.3 Limited access to computational resources.
Another barrier to ML reproducibility is the limited access to computational resources, which most recently has become evident in the case of transformer-based large language models (LLMs) [9, 79]. These transformer architectures need a vast amount of data and computational resources, to which most researchers have limited access. Estimates have calculated the costs to reproduce one model to be around \$1 million to \$3.2 million, which largely increases the difficulty of reproducing results [9]. Another study found that the needed computational resources are one of the most significant factors impacting reproducibility [62]. Especially ML models, which require computational clusters for training, are notably hard to reproduce.
4 Drivers for ML Reproducibility
In this section, we discuss drivers for ML reproducibility, which we subdivide into (i) technology-based drivers, (ii) procedural drivers, and (iii) drivers related to awareness and education.
4.1 Technology-based Drivers
4.1.1 Hosting services.
Utilizing hosting services offers an efficient way to share code, data, and ML model parameter settings, thus supporting the reproducibility of ML-driven research [71]. Examples of hosting services include the runtime environments of ML platforms. If the original author runs the ML experiment in such a runtime environment, e.g., Kaggle Kernels [^3] https://www.kaggle.com/, Google Colab [^4] https://colab.google/, or CodaLab [^5] https://codalab.org/, researchers attempting to reproduce the results should be able to execute the experiment within the same environment.
The main advantage of using a hosting service is that the provider takes care of the logistics of code hosting and distribution. However, the main drawbacks are the limits on data size and computational resources. Since these hosting services are run in the cloud, there are restrictions on how many resources a single user can utilize. The limit on resources varies between different hosting services and is sometimes tied to subscription levels. Because of these limits, hosting services may not be suitable for all research purposes, especially considering the compute-intensive nature of state-of-the-art ML models, such as LLMs. Furthermore, research has indicated that out-of-the-box reproducibility across multiple ML platforms is not yet guaranteed [29].
4.1.2 Virtualization.
Reproducing the environment and setup of any ML experiment requires the consideration of existing dependencies and software versions, and is usually a complex task itself. Virtualization can simplify this process by bundling the essential components of ML models and experiments, such as the environment and code, into a single package for being shared with other researchers. Thus, if the authors of a paper build the experiment in a virtual environment, issues associated with setup reproduction can be greatly reduced. However, the adoption of virtualization by researchers depends on its user-friendliness and the effort of integration into their current workflows [11]. Concerns about virtualization include its limitations in allowing researchers to build upon them in a scalable manner. Traditional virtual machines (VMs) emulate an entire operating system for setting up and running experiments. The use of containerization software like Docker [^6] https://www.docker.com/ has just recently become popular. Containers are more lightweight and flexible than VMs, making it easier to adapt environments for follow-up studies [11]. There are also designated platforms for computational research, such as Code Ocean [^7] https://codeocean.com/, that offer virtualization via so-called reproducible capsules. Their focus in particular is to simplify the virtualization process and allow researchers to focus on the research questions instead of the standardization of environments [15]. Additionally, there are many other tools, such as e.g., ReproZip [14], which is one of the recommended tools by the SIGMOD Reproducibility Committee to streamline reproducibility, and DetTrace [51], which aims to ensure completely deterministic computations.
4.1.3 Managing sources of randomness.
Many different sources of randomness during ML training lead to irreproducibility of ML models. Managing these sources of randomness, e.g., via random number seeds, deterministic algorithms, or other methods, could therefore greatly increase reproducibility. Random number seeds should be used and published to make ML experiments more reproducible and control a number of sources of the inherent nondeterminism. A seed is a first value used to initialize the pseudo-random number generator. When the same seed is used, the sequence of pseudo-random numbers generated is deterministic, meaning it will be the same every time the code is run. Additionally, ML models should be benchmarked and evaluated with multiple random number seeds, such that the variance can be reported and inform about the true performance of an ML model [63]. However, fixing the random number seeds can only eliminate algorithmic sources of randomness. Any differences due to implementation-level variance between low-level libraries could be overcome by implementing deterministic code, which has been demonstrated before [43]. Furthermore, hardware nondeterminism is also avoidable by using patching, i.e., replacing non-deterministic operations from hardware with deterministic ones, however, this will lead to a significant computing overhead [13]. To overcome the obstacle of inherent nondeterminism in Reinforcement Learning (RL), there are frameworks, which provide reproducible environments and are commonly used in RL research [12, 21, 78]. These frameworks act as benchmarking environments where different RL algorithms can be evaluated in a common environment. Additionally, different papers have outlined the implementation of deterministic RL algorithms, by controlling all the sources of non-determinism [50].
4.1.4 Privacy-preserving technologies.
Privacy-preserving technologies support reproducibility, as they enable the collaborative training of ML models without sharing private or sensitive data. The main benefit of this is, that ML models can make use of larger, and more diverse data, thus helping to decrease bias and leading to more reproducible ML models. The main aim of Privacy-Preserving Machine Learning (PPML) is to facilitate the use of privacy-sensitive data to create better ML models, and, to allow data owners to collaboratively train ML models on private data. In that regard, PPML has several requirements. First, protecting the confidentiality of the training data. Second, preventing the leakage of sensitive information from ML model parameters and outputs, i.e., to hinder the re-identification of individuals. Third, achieving the listed security aspects while still preserving the utility of the ML model [76]. To achieve this, a number of different techniques are being used and developed, such as differential privacy, fully homomorphic encryption, secure multi-party computation, and federated learning [19]. These techniques are implemented in software libraries such as TensorFlow Privacy, PySyft, ML Privacy Meter, CryptFlow, or Crypten [7]. Furthermore, data can be made anonymous by removing identifiable personal information. However, if too much data is removed, the ML models may perform poorly. If not enough data is removed, it may still be possible to re-identify individuals by combining many different non-unique features [76]. An alternative approach to PPML is to generate synthetic data that captures the same information as the original data. A robust technique for creating such datasets can produce readily available datasets of nearly any size, as demonstrated in Biomedical fields. This approach has led to the development of Synthea, a software package designed to generate synthetic patients and electronic health care records [73].
4.1.5 Tools and platforms.
There are many tools and platforms that assist in the implementation and management of ML model and ML-based applications ranging from . Some of them provide functionalities to increase ML reproducibility. A recent study has evaluated 19 ML tools to gain insight into their concepts constituting reproducibility support [61]. As a result, five main pillars of ML reproducibility in tools and platforms were identified: (i) code versioning, (ii) data access, (iii) data versioning, (iv) experiment logging, and (v) pipeline creation. Most of these pillars are associated with managing and keeping track of different artifacts created during phases of the ML lifecycle (i.e., design, development, and deployment) as for instance, datasets, labels, code, logs, environment dependencies, random number seeds or hyperparameters [67]. Each of these artifacts influences the final results of the ML model. Consequently, most tools aim to collect, store, and manage these artifacts, ensuring researchers can access and use them during reproduction attempts. Notable are also various tools and platforms for experiment tracking [67]:
  • DVC [^8] https://dvc.org/ is a version control system for ML projects with a command-line interface similar to Git [^9] https://git-scm.com/. It integrates with Git, supports cloud storage, and handles large dataset versioning. DVC ensures full code and data provenance by enabling experiment tracking.
  • MLflow [^10] https://mlflow.org/ is an open-source tool for supporting ML experiment tracking, ML model deployment, and centralized model storage. Additionally, it provides an easy-to-use user Web dashboard.
  • RO-Crate is a specification, implemented by a number of tools, aimed at aggregating and describing research data and metadata [69]. Although not specifically designed for ML, RO-Crate can aggregate and represent any resource, making it applicable for managing ML artifacts as well.
  • dToolAI [31] collects and packages ML models together with supplemental information, such as hyperparameters, appropriate metadata, and persistent URIs for model training data. In contrast to the other tools, dToolAI is specifically tailored for deep learning models.
AutoML platforms, such as H2O Driverless AI [^11] https://h2o.ai, Google Cloud AutoML [^12] https://cloud.google.com/automl, DataRobot [^13] https://www.datarobot.com/platform) are a novel subcategory of ML tools that aim to aid with every aspect of the ML lifecycle, from data aggregation to model deployment. AutoML tools could thus facilitate more standardized ML models and also take care of tasks like hyperparameter optimization. Furthermore, many of them also provide a lot of the features identified as the main pillars of reproducibility in ML tools. It is however questionable how practical these tools are for research, since they often act as black-boxes without the need to apply programming expertise. Sometimes, e.g., in the case of H2O Driverless AI, the features go a step further and promise to be able to prevent problems such as overfitting or target leakage. In the case of leakage, this is done by checking for a strong correlation between a feature and the target and then taking action, e.g., warning the user or automatically handling it. This is, however, a very simple solution to the problem and does not address the more complex cases of data leakage that are often present in research, e.g., temporal leakage. [^14] https://h2o.ai/wiki/target-leakage//
4.2 Procedural Drivers
4.2.1 Standardized datasets and evaluation.
Due to a lack of shared datasets, currently, many researchers in ML-driven research - most notably in Biomedical fields - use individually acquired data [45]. The collection of such data is a time-consuming task and bears a significant risk of causing reproducibility issues, e.g., bias or data leakage. Often, the data population of these datasets is not very large and, thus, might not translate well to new contexts. Creating shared and standardized datasets can, therefore, (i) save researchers time in acquiring new data, (ii) facilitate the collaborative and independent maintenance and verification of data to minimize methodological errors, and (iii) support transferability and generalizability, through the use of multi-institutional data [45]. In addition to the standardization of datasets, data cards [60] provide a consistent and comparable framework for reporting essential aspects of ML datasets. This includes information e.g., about access restrictions, risks and limitations associated with the usage of the dataset, or any preprocessing steps, amongst many other contents, which are needed for responsible ML development.
Another issue is the lack of standardized evaluation, which leads to reported performances of ML models often being overly optimistic [18]. To ensure the statistical significance of ML model evaluations, it is crucial to report performance as an aggregate of results obtained from multiple random runs with different random number seeds [16]. Furthermore, ML models should, if possible, be tested and evaluated on multiple different datasets [20]. This underscores the need for standardized evaluation methods, which can be supported by checklists or tools to prevent errors in this critical aspect of ML research. For this, similarly to data cards, model cards [46] are aimed to standardize the evaluation and reporting of the performance of ML models for a variety of use cases. Model cards should inform users about the possible applications of the ML model and its limitations. For the creation of the model cards, a toolkit is also being developed by Google [^15] https://research.google/blog/introducing-the-model-card-toolkit-for-easier-model-transparency-reporting/. In RL, the creation of standardized evaluation pipelines is continually being researched to enable reproducible benchmarking of different RL algorithms [40].
4.2.2 Guidelines and checklists.
There are many guidelines and checklists that outline best practices for increasing the reproducibility of ML. The guidelines are often aimed at specific parts of the ML workflow. The FAIR guiding principles [^16] https://www.go-fair.org/fair-principles/, e.g., aim to improve the management and stewardship of scientific data by making scientific data (i) findable, (ii) accessible, (iii) interoperable, and (iv) reusable. Other guidelines promote the transparency and openness of scientific reporting in general, such as the TOP guidelines [^17] https://www.cos.io/initiatives/top-guidelines, which target journals. Similarly, checklists provide a simple framework for ensuring certain criteria are met. Checklists have been applied effectively in the past, e.g., in safety-critical systems, where they were used as early as in 1935 to complete pre-flight checks in Boeing airplanes. A promising example is the ML checklist proposed in [57]. It has been mentioned during the NeurIPS 2019 reproducibility challenge, and has been suggested as best practice by researchers of different fields, e.g., in chemistry [6]. The checklist requests information about (1) the models and algorithms being used, (2) theoretical claims in the research article, (3) data, (4) code, and (5) the ML experiment(s). However, one drawback of reproducibility checklists when used for academic conferences and journals is the additional workload they impose on reviewers. To mitigate this workload, one option would be to leverage LLMs to assist the review process [44]. Finally, numerous guidelines and checklists for ML reproducibility have been recommended in various research fields [6]. Especially in Biomedical fields, we have seen a considerable adoption of guidelines and checklists, such as the TRIPOD statement [17], the CLAIM checklist [47], the ROBUST-ML checklist [2] or PROBAST [75].
4.2.3 Model info sheets and model cards.
Model cards [46] are documentation sheets that provide information about ML models, including their intended use, potential limitations, and ethical considerations. They aim to enhance transparency in AI, by detailing aspects such as data used for training, performance metrics and evaluation methodologies, and possible biases. Model cards help users to understand and evaluate ML models more comprehensively, such that they are not being deployed in contexts other than the intended use, and thus to increase reproducibility. Similarly, model info sheets also provide documentation about ML models, but are specifically designed for the detection and prevention of data leakage in ML models [39]. The model info sheets are published alongside research to enable other researchers to quickly verify the validity of the data used to train ML models. Model info sheets require authors to answer detailed questions about the data and corresponding train/test splits, targeting various types of data leakage [39]. Despite that, they have two main drawbacks: first, verifying the correctness of info sheets only works after reproducing the results; second, completing these sheets requires a certain level of expertise in ML.
In general, model cards and model info sheets represent a promising, low-effort driver for ML reproducibility. They are especially useful in handling some of the methodological issues associated with ML models being developed by non-experts [39].
4.3 Awareness and Education
Awareness of reproducibility issues and available training/education to support reproducibility can be a very strong, contributing factor to improve the implemented practice (e.g., [74]).
4.3.1 Publication policies and initiatives.
To enhance awareness and establish a minimum of reproducibility standards, the policies of scientific journals are considered an influence factor. A number of journals already mandate data and/or code availability for publication [57, 56, 30]. However, to address issues such as result manipulation, more extensive journal participation is needed to for instance introduce preregistration, where researchers register their research intentions for future publication. This approach ensures credibility by separating the research plan from experimental outcomes [70, 52], thereby reducing spin practices, HARKing, and p-hacking [29]. ACM TORS (Transactions on Recommender Systems) journal exemplifies this by allowing preregistration and publishing “reproducibility papers” dedicated to reproduction studies and enhancing reproducibility tools. Apart from that, various initiatives have been launched to raise awareness of reproducibility issues. A few examples are provided below:
  • The ReScience journal publishes peer-reviewed papers discussing attempts to reproduce original publications. These reproductions are published on GitHub [^18] https://github.com/, being available to other researchers [64].
  • PapersWithCode.com [^19] https://paperswithcode.com/ is a resource for (1) ML papers, accompanied by the code, (2) datasets, and (3) ML methods. The ML papers include a link to a repository, which features the code and other artifacts for reproducing the results.
  • Reproducibility challenges, where several researchers try to reproduce many recent publications in parallel, are being held frequently. These challenges allow for an analysis of the success rate of reproduction and can be used to evaluate progress over multiple years [57]. Additionally, conferences such as the European Conference on Information Retrieval (ECIR), provides special reproducibility tracks, in which researchers are encouraged to reproduce existing papers and build upon their results (e.g., [42, 48]).
  • The ACM has put together a new emerging interest group on reproducibility [^20] https://reproducibility.acm.org/. The main goals are to (1) contribute to the development of reproducibility standards, practices and policies, (2) promote the development and evaluation of tools, techniques, and methodologies and (3) encourage best practices.
  • ReproducedPapers.org is another online repository fostering reproductions. It further, focuses on education by incorporating a reproduction project into a Master’s level ML course at TU Delft [77].
5 Mapping Drivers to Barriers
In this section, we will map the drivers of reproducibility to the barriers in the form of a Drivers-Barriers Matrix. This will be based on the definition and categorization of reproducibility as a foundation (Section 2), and the identification of the major barriers (Section 3) and drivers (Section 4) of ML reproducibility. The resulting Drivers-Barriers-Matrix is depicted in Figure 2 and categorizes the barriers into the four different types of ML reproducibility, i.e., R1 Description, R2 Code, R3 Data, R4 Experiment [25], and drivers into technology-driven, procedural, and related to awareness and education.
Our Drivers-Barriers-Matrix shows that there are often multiple drivers for the same barrier. Consequently, there are also several possible solutions for a barrier or different aspects of a barrier. The mapping allows to quickly assess which drivers address the different barriers and which barriers have a higher or lower number of drivers associated with them. It underlines the need for context-dependent approaches instead of "one-size-fits-all" solutions, as the proper selection of a suitable driver depends on the specific conditions and existing barriers relevant to any ML application. Intersections between drivers and barriers will be described in more detail below.
5.0.1 R1 Description.
Completeness and quality of reporting, as well as spin practices and publication bias, present the major barriers associated with R1 Description. These are characterized as missing information in reports and overinflated results that hinder reproducibility.
The major drivers for completeness and report quality are guidelines and checklists. Guidelines provide best practices to adopt in order to achieve reproducible ML research. Furthermore, many checklists exist that comprehensively state the different pitfalls and provide information on how they can be avoided. Researchers can use them to ensure their research meets the desired standards. Furthermore, some of these checklists and guidelines are enforced by journals, such that research will only be published if certain criteria are met.
In comparison, spin practices are not as easily identifiable. In this case, the discussion within the research community centers around removing the incentives for inflating research results. A particularly effective driver for this is preregistration as an example for publication policies and initiatives, where researchers submit research proposals before conducting an experiment. If accepted, the research will be published, regardless of the outcome, thereby minimizing spin practices.
Figure 2: Drivers-Barriers-Matrix. We map the 9 drivers to the 9 barriers identified in this paper. The colored boxes show that a specific driver is applicable to a specific barrier.
5.0.2 R2 Code.
Code sharing is essential to reproducibility, which makes limited access to code a significant barrier in the field. However, it is often neglected as the process is not trivial. To make shared code useful to the scientific community, it is necessary to share, in addition to the source code, the information about the entire software setup and dependencies, including software versions and hardware configurations. To assist with this, researchers may consider running their code in hosting services or virtualization environment, which we identified as drivers for code sharing. Both have similar advantages, i.e., they can easily be shared and made public for other researchers to use. As a consequence, it will give reproducers immediate access to code, including the complete configuration setup, such as dependencies and versions. Hosting services are the quicker and easier way of achieving this; however, they may be subject to different resource limits. Virtualization (e.g., VMs or containers) is more difficult to set up but offers more flexibility and is not externally (e.g., by a provider) restricted in capabilities and resources. Furthermore, tools and platforms can be drivers for reproducibility. A lot of ML tools provide capabilities for code versioning or other features, which are key to reproducibility. One example is dToolAI [31], which automatically logs the supplemental information of the code, i.e., important metadata, hyperparameters, and more, which are essential to the reproducibility of type R2 Code.
5.0.3 R3 Data.
Data-related barriers are a severe obstacle to ML reproducibility due to the research fields’ data-driven nature, where limited access to data forms a major challenge. Privacy concerns are among the crucial arguments that cause hesitation in sharing data. The need for data privacy is evident, especially in the Biomedical fields, which deal with patients’ electronic health records. Nevertheless, it increases reproducibility issues in ML-based science and, thus, delays technological progress within these domains. However, there are several approaches that aim to meet the requirements of sharing sensitive data: Privacy-preserving technologies allow reproducers to train ML models on private data without actually possessing the data. This way, reproduction becomes possible without violating potential privacy regulations. Other than that, the use of standardized datasets and evaluation can support issues in regard to dataset meta-information, including the specification of train-test splits and data provenance. Once again, tools and platforms can assist with data versioning, and numerous guidelines and checklists have been proposed to address the provenance of data. These guidelines and checklists are designed to help researchers to avoid common pitfalls. Current initiatives are supported by journals that more frequently require data to be shared as part of a publication.
Concerning methodological errors associated with the data, data leakage is a major issue, which can, for instance, be mitigated using standardized datasets and evaluation. Other drivers to solve data leakage are model info sheets and model cards, which are provided as supplemental information to a published dataset. Even though there are some limitations to model info sheets, they are capable of detecting all types of data leakage.
Bias is another methodological error, leading to irreproducible results. This is because the biased data usually does not generalize well to problems outside the experiment setup. Biases have been an important source of concern, e.g., in Biomedical fields. Effects thereof can again be minimized using standardized datasets and evaluation or specific guidelines and checklists, e.g., ROBUST-ML [2].
5.0.4 R4 Experiment.
If an ML experiment is shared entirely and code and data are available, i.e., reproducibility type R4 Experiment, there are still three barriers, which can lead to irreproducible results. Inherent non-determinism arises from the different sources of randomness in ML, and makes it difficult to achieve repeatable results on the same machine. There are, however, methods to manage the sources of randomness, such as fixed random seeds and deterministic implementations, while comprehensively mitigating all sources of randomness is still a very challenging endeavor.
Another barrier is described as environmental differences, which has two main issues associated with it, i.e., software differences and hardware differences. Both types of differences can be avoided by using either hosting services or virtualization; constraints can be assumed to be similar to the barrier of limited access to code.
Limited access to computational resources constitutes another barrier to ML reproducibility identified in this work. The issue is particularly problematic for research on large language models because of their need for extensive computational resources in training and reproduction. Hosting services offer a solution, providing access to pre-trained models and allowing researchers to directly access and run respective models on-site.
6 Conclusion
In this paper, we examined the barriers and drivers associated with the four types of ML reproducibility as outlined by Gundersen et al. (Description, Data, Code, and Experiment) [25], specifically in the contexts of Computer Science and Biomedical research. We then synthesized this into a Drivers-Barriers-Matrix to summarize and illustrate which drivers are feasible solutions for the various barriers. We observe that the barriers to machine learning reproducibility can be addressed through three primary drivers: technology-driven solutions, procedural improvements, and enhanced awareness and education. It is important to highlight that, in theory, awareness and education can complement the other drivers and serve as a foundational step toward overcoming any reproducibility-related challenge. Further, the rise of AutoML tools for ML development and ML tasks performed by domain experts (potentially) not having in-depth computer or data science knowledge could become another barrier for reproducibility. However, we are confident that our findings will also help in this regard and that the topic likely requires collaborative efforts over the coming years.
We hope that this provides practical guidance and orientation for ML researchers, and clarifies the current state of play. Of course, in such a dynamic and fast-paced research area, this only opens a series of further questions and avenues for exploration. We recommend further investigation of the various issues and potential solutions laid out here. We would also encourage further investigation into the potential role of platforms [29] or foundation models [34] in further exacerbating or alleviating these issues.
6.0.1 Acknowledgements.
This research is supported by the Horizon Europe project TIER2 (GA: 101094817), and the FFG COMET program.